### v1

# Parameter Invariance Analysis of Moment Equations Using Dulmage-Mendelsohn Decomposition

Abstract
Living organisms maintain stable functioning amid environmental fluctuations through homeostasis, a mechanism that preserves a system’s behavior despite changes in environmental conditions. To elucidate homeostasis in stochastic biochemical reactions, theoretical tools for assessing population-level invariance under parameter perturbations are crucial. In this paper, we propose a systematic method for identifying the stationary moments that remain invariant under parameter perturbations by leveraging the structural properties of the stationary moment equations. A key step in this development is addressing the underdetermined nature of moment equations, which has traditionally made it difficult to characterize how stationary moments depend on system parameters. To overcome this, we utilize the Dulmage-Mendelsohn (DM) decomposition of the coefficient matrix to extract welldetermined subequations and reveal their hierarchical structure. Leveraging this structure, we identify stationary moments whose partial derivatives with respect to parameters are structurally zero, facilitating the exploration of fundamental constraints that govern homeostatic behavior in stochastic biochemical systems.
I Introduction
Homeostasis is the property by which organisms adapt to changes in external environments, enabling survival in highly uncertain conditions. Experimental studies reported that homeostasis also functions at the cellular scale within chemical reaction networks[1, 2, 3]. However, many underlying principles of homeostasis remain unclear, highlighting the need for model-based analysis. Mathematically, homeostasis can be described as the ability of the reaction network to maintain its state variables invariant to perturbations in parameters such as reaction rates, which may be affected by environmental factors such as temperature and osmorarity [4, 1, 5]. To date, many efforts have been made to analyze the invariance based on deterministic models, specifically the rate equations. These include approaches that utilize reaction stoichiometry to identify subnetworks attaining perfect adaptation [6, 7, 8] and those employing sensitivity analysis based on the Jacobian matrix [9, 10, 11]. However, since intracellular reactions are inherently stochastic due to the low copy number of molecules, developing methods based on stochastic models is desirable for a more accurate analysis that better aligns with experimental observations.
The governing equation for stochastic chemical reactions is the chemical master equation (CME), an infinite-dimensional linear differential equation for a discrete-state Markov process that poses challenges for sensitivity and invariance analysis [12, 13]. To address this, various numerical methods were proposed to evaluate the sensitivity of the density distributions of the copy numbers using the Fisher information matrix (FIM), employing Monte Carlo simulations or model approximations [14, 15]. Meanwhile, moment-based approaches are amenable to analytic or algebraic characterization of stationary moments based on the moment equation [16, 17] and enabled rigorous analysis of invariant moments for a certain class of biomolecular networks. A notable example is the antithetic integral feedback (AIF) network, where the stationary mean copy numbers can be explicitly given in terms of biochemical parameters, allowing the theoretically guaranteed design of perfect adaptation for stochastic chemical reactions [18]. However, no existing methods achieve both (i) a mathematically rigorous analysis of parameter invariance and (ii) general applicability to arbitrary chemical reaction networks for stochastic systems. This motivates the development of a systematic approach for analyzing parameter invariance in stochastic chemical systems that can rigorously address these limitations.
In this study, we propose a systematic method for analyzing the parameter invariance of stationary moments in general chemical reaction networks by leveraging the structural information of the moment equations. We utilize the Dulmage-Mendelsohn (DM) decomposition [19, 20, 21], a graph-based matrix operation that extracts a welldetermined subequations from a system of linear equations, to identify determinable variables in the highly underdetermined stationary moment equations and the dependency structure among the variables. By linking this structure to unreachability analysis in directed graphs, we establish a general algorithm to systematically identify stationary moments whose partial derivatives with respect to parameters are structurally zero. The proposed method contributes to a broader understanding of homeostasis while explicitly accounting for the stochastic nature of biochemical reactions.
Notations: The symbol $\mathbb{P}(\cdot)$ is a probability mass function, and $\mathbb{E}[f(\cdot)]$ is the moment of the mass function defined by $\sum_{\bm{x}}f(\bm{x})\mathbb{P}(\bm{x})$. The expression $\bm{x}^{\bm{\alpha}}$ with $\bm{x}=[x_{1},x_{2},\ldots,x_{n}]^{\top}$ and $\bm{\alpha}=[\alpha_{1},\alpha_{2},\ldots,\alpha_{n}]^{\top}$ denotes multivariate monomial of $x_{1}^{\alpha_{1}}x_{2}^{\alpha_{2}}\dots x_{n}^{\alpha_{n}}$.
II Problem Formulation
In this section, we introduce a model describing the stochastic dynamics of chemical reactions and formulate the parameter invariance analysis problem.
II-A Chemical Master Equations
We consider chemical reaction networks $\Gamma$ consisting of $n$ molecular species $X_{i}\>(i=1,2,\dots,n)$ and $m$ reaction species $R_{j}\>(j=1,2,\dots,m)$. In the modeling of stochastic chemical reactions, the state of the reaction networks $\Gamma$ is characterized by the copy number of molecular species $X_{i}$, and its evolution is modeled by the Markov process governed by stochastic increase and decrease of the copy numbers by reactions $R_{j}$.
To formally define this stochastic process, let $x_{i}$ denote the copy number of the molecular species $X_{i}$, and define the state of the reaction network $\Gamma$ at time $t$ by $\bm{x}(t)=[x_{1},x_{2},\dots,x_{n}]^{\top}\in\mathbb{N}_{0}^{n}$. Reaction species $R_{j}$ is characterized by a stoichiometry vector $\bm{s}_{j}\in\mathbb{Z}^{n}$, which describes the change in $\bm{x}$ upon the occurrence of the reaction, and a propensity function $w_{j}(\bm{x})$ representing the probability of its occurrence per unit time. The propensity function $w_{j}(\bm{x})$ is defined for each reaction type by
$\!w_{j}(\bm{x}):=\begin{cases}\theta_{j}&\text{zero-th order reaction}\\ \theta_{j}x_{i}&\text{monomolecular reaction}\\ \theta_{j}x_{i}x_{l}&\text{bimolecular of $X_{i}$ and $X_{\ell}$}\\ {\theta_{j}x_{i}(x_{i}-1)}/2&\text{bimolecular of $X_{i}$'s}\\ \end{cases}$
with the rate parameters $\theta_{j}\in\mathbb{R}_{+}$, assuming that $R_{j}$ is one of the elementary reactions [5].
The stochastic chemical reactions are then modeled as the evolution of the copy number distribution of the state $\bm{x}$ at time $t$ denoted by $\mathbb{P}\left(\bm{x}|\bm{\theta},t,\bm{x}_{0}\right)$, for a given initial state $\bm{x}_{0}$, i.e., $\bm{x}_{0}:=\bm{x}(0)$. The governing equation of the dynamics is called the Chemical Master Equation (CME) [12, 13], or Kolmogorov’s forward equation. Omitting $t$ and $\bm{x}_{0}$ in $\mathbb{P}\left(\bm{x}|\bm{\theta},t,\bm{x}_{0}\right)$ to avoid notational clutter, the CME can be expressed as
$\!\!\!\!\!\frac{d\mathbb{P}(\bm{x}|\bm{\theta})}{dt}\!=\!$ $\sum_{j=1}^{m}\{w_{j}(\bm{x}\!-\!\bm{s}_{j})\mathbb{P}(\bm{x}\!-\!\bm{s}_{j}|\bm{\theta})\!-\!w_{j}(\bm{x})\mathbb{P}(\bm{x}|\bm{\theta})\}\!\!\!\!$ (1)
for each $\bm{x}\in\mathbb{N}_{0}^{n}$, where $\bm{\theta}:=[\theta_{1},\theta_{2},\ldots,\theta_{m}]^{\top}\in\mathbb{R}_{+}^{m}$. It should be noticed that the CME is a linear but an infinite dimensional ordinary differential equation (ODE), for which analytic solutions are rarely available.
II-B Problem statement
In the design and analysis of stochastic reaction networks, one is often interested in understanding how the copy number distributions $\mathbb{P}(\bm{x}|\bm{\theta})$ depend on the rate parameters $\bm{\theta}$, which can be influenced by environmental changes such as temperature and osmolarity[5]. In this work, we examine these dependencies through the moments of the copy number distribution $\mathbb{E}[\bm{x}^{\alpha}|\bm{\theta}]$, which capture essential statistics such as the mean and the variance. More specifically, our focus is on the stationary moments, and we investigate whether certain stationary moments remain invariant under changes in the rate parameters $\bm{\theta}$, as such invariance is closely related to the homeostatic behavior of biomolecular systems.
To formally define the problem, we introduce the following assumption.
Assumption 1
For any rate parameters $\bm{\theta}$ and initial state $\bm{x}_{0}$, (i) the stationary solution of the CME (1) exists, and (ii) its associated Markov process is non-explosive. Moreover, (iii) all moments of the stationary distributions are finite.
This assumption guarantees the existence of the moments of the stationary copy number distribution, or the stationary moments. In what follows, we denote the stationary moments by $\mu_{\bm{\alpha}}(\bm{\theta}):=\mathbb{E}[\bm{x}^{\bm{\alpha}}|\bm{\theta}]$, where $\bm{\alpha}\in\mathbb{N}_{0}^{m}$ is a multi-index specifying the moment (see Notations in Section I). Then, the problem is formally stated as follows.
Problem 1
For a given chemical reaction network $\Gamma$ and a reaction species $R_{k}$ to be perturbed, assume that Assumption 1 holds. Define the variation of the parameter associated with the reaction species $R_{k}$ around $\bm{\theta}$ by
$\Delta_{k}(\bm{\theta}):=\{\bm{\theta}+\phi\bm{\delta}_{k}\mid\phi\in\mathbb{R},\bm{\theta}+\phi\bm{\delta}_{k}\in\mathbb{R}_{+}^{m}\},$ (2)
where $\bm{\delta}_{k}$ is the $k$-th standard basis of $\mathbb{R}^{m}$ with the $k$-th entry being 1 and the others being 0. Under this parameter variation, find a set of invariant stationary moments
$\mathcal{M}_{k}:=\left\{\bm{\alpha}\ \middle|\ \begin{aligned} &{\mu}_{\bm{\alpha}}(\bm{\theta}+\delta\bm{\theta})={\mu}_{\bm{\alpha}}(\bm{\theta})\\ &\forall\bm{\theta}\in\mathbb{R}^{m}_{+},\ \forall\delta\bm{\theta}\in\Delta_{k}(\bm{\theta})\end{aligned}\right\}.$ (3)
III Analysis of Invariant Moments
In this section, we present a systematic method for identifying the invariant stationary moments $\mathcal{M}_{k}$. Our approach analyzes the sensitivity of stationary moments with respect to the parameters and derives an algebraic equation that characterizes this sensitivity. We then develop an algorithm that systematically explores the structural properties of this equation to identify the invariant stationary moments.
III-A Overview of the proposed approach
Let $\mu^{(1)}_{\bm{\alpha}}(\bm{\theta})$ denote the partial derivative of the stationary moment with respect to $\theta_{k}$ defined as
$\mu^{(1)}_{\bm{\alpha}}(\bm{\theta}):=\frac{\partial\mu_{\bm{\alpha}}}{\partial\theta_{k}}\bigg{|}_{\bm{\theta}}.$ (4)
By definition, the stationary moment $\mu_{\bm{\alpha}}(\bm{\theta})$ is invariant under the change in the parameter $\theta_{k}$, i.e., $\bm{\alpha}\in\mathcal{M}_{k}$ holds, if and only if its derivative satisfies $\mu^{(1)}_{\bm{\alpha}}(\bm{\theta})=0$ for all $\bm{\theta}$. In other words, our problem reduces to identifying the moments that remain invariant due to the structural properties of the system, independent of the specific parameter values $\bm{\theta}$.
To explore such moments, we consider the stationary moment equation. Specifically, we first obtain the linear equation for the stationary distribution by setting the left-hand side of eq. (1) to zero. Next, we multiply these linear equation by $\bm{x}^{\bm{\alpha}}$ and sum over all $\bm{x}\in\mathbb{N}^{n}_{0}$ to derive the moment equation [16]
$\sum_{\bm{\beta}}\sum_{j=1}^{m}\theta_{j}(a^{j}_{\bm{\alpha},\bm{\beta}}\mu_{\bm{\beta}}(\bm{\theta})+b^{j}_{\bm{\alpha}})=0,$ (5)
where $a^{j}_{\bm{\alpha},\bm{\beta}}$ and $b^{j}_{\bm{\alpha}}$ are coefficients determined by the stoichiometry vectors $\bm{s}_{j}$ and propensity functions $w_{j}(\bm{x})$. Equation (5) is a set of linear equations that the stationary moment $\mu_{\bm{\alpha}}(\bm{\theta})$ must satisfy, and the coefficient $a^{j}_{\bm{\alpha},\bm{\beta}}$ represents the interaction between the stationary moments $\mu_{\bm{\alpha}}$ and $\mu_{\bm{\beta}}$ through reaction $R_{j}$, while $b^{j}_{\bm{\alpha}}$ is a constant term arising from zero-th order reactions or conserved quantities of molecular copy number. Taking the derivative of eq. (5) with respect to ${\theta}_{k}$, we have
$\sum_{\bm{\beta}}\left\{a^{k}_{\bm{\alpha},\bm{\beta}}\mu_{\bm{\beta}}(\bm{\theta})+b^{k}_{\bm{\alpha}}+\sum_{j=1}^{m}\theta_{j}a^{j}_{\bm{\alpha},\bm{\beta}}\mu^{(1)}_{\bm{\beta}}(\bm{\theta})\right\}=0,$ (6)
which is a linear equation with respect to the moments and their partial derivatives with respect to $\theta_{k}$. In typical cases where the reaction network $\Gamma$ includes bimolecular reactions, the stationary moment equation is underdetermined since each moment equation depends on higher order moments. This dependence leads to an analytically intractable infinite dimensional equation, and thus, it is not straightforward to analyze the parameter dependence of the moments.
To address this challenge, we present an approach that utilizes the structural information of these algebraic equations to identify the invariant moments satisfying $\mu_{\bm{\alpha}}^{(1)}(\bm{\theta})=0$ for all $\bm{\theta}\in\mathbb{R}_{+}^{m}$. To this end, we first replace the variables (moments) in eq. (5) and eq. (6) with independent variables and define the following linear equations
$\!\!\begin{bmatrix}\sum_{j=1}^{m}\theta_{j}A_{j}&A_{k}\\ O&\sum_{j=1}^{m}\theta_{j}A_{j}\end{bmatrix}\begin{bmatrix}\bm{\nu}^{(1)}\\ \bm{\nu}\end{bmatrix}=-\begin{bmatrix}\bm{b}_{k}\\ \sum_{j=1}^{m}\theta_{j}\bm{b}_{j}\end{bmatrix},$ (7)
where $A_{j}\in\mathbb{R}^{M\times N}(j=1,2,\ldots,m)$ and $\bm{b}_{j}\in\mathbb{R}^{M}(j=1,2,\ldots,m)$ are matrices and vectors that contain the coefficients $a_{\bm{\alpha},\bm{\beta}}^{j}$ and $b_{\bm{\alpha}}^{j}$, respectively, and $\bm{\nu}$ and $\bm{\nu}^{(1)}$ are the vectors of independent variables associated with the stationary moments and their derivatives. For simplicity, we make the following assumptions that eliminate redundant equations and unnecessary variables.
Assumption 2
For all $\bm{\theta}\in\mathbb{R}^{m}_{+}$, the matrix $\sum_{j=1}^{m}\theta_{j}A_{j}$ in eq. (7) is full row rank, and at least one entry of for each column of $\sum_{j=1}^{m}\theta_{j}A_{j}$ is non-zero.
The full row rank condition implies that the moment equation does not have redundant equations, which can be systematically eliminated by utilizing the basis vectors common to the left null spaces of the augmented coefficient matrices $[A_{j}\ \bm{b}_{j}]\>(j=1,2,\dots,m)$. The latter half of the assumption ensures that the vector $[\bm{\nu}^{(1)},\bm{\nu}]^{\top}$ does not contain unused variables.
Next, to identify the moments satisfying $\bm{\mu}_{\bm{\alpha}}^{(1)}(\bm{\theta})=0$ for all $\bm{\theta}$, we focus on the fact that the arrangement of non-zero elements in the matrices $A_{j}$ and the vectors $\bm{b}_{j}$ are independent of $\bm{\theta}$. In particular, we utilize the zero and non-zero structure of eq. (7) and explore the entries of $\bm{\nu}^{(1)}$ that are forced to be zero due to the structural constraint. For this purpose, we define an operator $\pi:\mathbb{R}^{M\times N}\cup\mathbb{R}^{M}\rightarrow\{0,1\}^{M\times N}\cup\{0,1\}^{M}$ that maps a matrix or vector to a binary matrix or vector encoding the zero and non-zero structure, respectively. Specifically, for a matrix $X\in\mathbb{R}^{M\times N}$ and a vector $\bm{x}\in\mathbb{R}^{M}$,
$\!\!\!(\pi(X))_{p,q}\!:=\!\!\begin{cases}\!1\!\!\!\!&\!\text{if $(X)_{p,q}\!\neq\!0$}\\ \!0\!\!\!\!&\!\text{otherwise}\end{cases}\!,(\pi(\bm{x}))_{p}\!:=\!\!\begin{cases}\!1\!\!\!&\!\!\text{if $(\bm{x})_{p}\!\neq\!0$}\\ \!0\!\!\!&\!\!\text{otherwise}\end{cases}.\!\!$
Our approach is then to seek invariant moments for all possible equations that share the same zero and non-zero structure as the stationary moment equations. The following Proposition is instrumental for the analysis of invariant moments.
Proposition 1
Consider linear equations
$\Phi\bm{\nu}_{\Phi}=\bm{b}_{\Phi},$ (8)
where $\Phi\in\mathbb{R}^{2M\times 2N}$ and $\bm{b}_{\Phi}\in\mathbb{R}^{2M}$ are the coefficient matrix and vector that share the same non-zero structure as the coefficient matrix and vector eq. (7), i.e.,
$\Phi:=\begin{bmatrix}\tilde{A}&\tilde{A}_{k}\\ O&\tilde{A}\end{bmatrix},\bm{\nu}_{\Phi}$ $:=\begin{bmatrix}\tilde{\bm{\nu}}^{(1)}\\ \tilde{\bm{\nu}}\end{bmatrix},\bm{b}_{\Phi}:=\begin{bmatrix}\tilde{\bm{b}}_{k}\\ \tilde{\bm{b}}\end{bmatrix}$ (9)
with $\tilde{\bm{\nu}}^{(1)}$ and $\tilde{\bm{\nu}}$ being vectors with the same size as ${\bm{\nu}}^{(1)}$ and ${\bm{\nu}}$, respectively, $\tilde{A}_{j}\in\{X\in\mathbb{R}^{M\times N}\mid\pi(X)=\pi({A}_{j})\},\tilde{A}\in\{X\in\mathbb{R}^{M\times N}\mid\pi(X)=\bigcup_{j=1}^{m}\pi({A}_{j}),\text{rank}(X)=M\},\tilde{\bm{b}}_{j}\in\{\bm{x}\in\mathbb{R}^{M}\mid\pi(\bm{x})=\pi(\bm{b}_{j})\}$ and $\tilde{\bm{b}}\in\{\bm{x}\in\mathbb{R}^{M}\mid\pi(\bm{x})=\bigcup_{j=1}^{m}\pi(\bm{b}_{j})\}$. The union of the binary matrices and vectors are defined so that the entry becomes 1 if at least one of the matrices and vectors has a non-zero entry in the corresponding position, and 0 otherwise, respectively. If the $i$-th entry of the solution $\bm{\nu}_{\Phi}$ is zero for every solution of eq. (8), the $i$-th entry of $[\bm{\nu}^{(1)},\bm{\nu}]^{\top}$ is also zero for all $\theta_{j}\in\mathbb{R}_{+}\ (j=1,2,\ldots,m)$ and for every solution of eq. (7).
This proposition implies that the partial derivative (4) is zero if the corresponding entry of $\bm{\nu}^{(1)}$ is zero for all possible equations having the same non-zero structure as eq. (7). In the next subsections, we show an approach that systematically identifies such entries of $\bm{\nu}^{(1)}$ of eq. (7).
III-B Dulmage-Mendelsohn decomposition
We analyze the solutions of the linear equations (8) by systematically permuting the rows and columns of the coefficient matrix $\Phi$ to make triangular blocks corresponding to welldetermined and underdetermined subequations. The entries of the solution $\bm{\nu}_{\Phi}$ that are structurally constrained to be zero are then obtained from the welldetermined subequations. For this purpose, we utilize the Dulmage-Mendelsohn (DM) decomposition [19, 20, 21], originally introduced as a partitioning algorithm for bipartite graphs. The following lemma shows the existence and the uniqueness of such block-triangular decomposition of the matrix $\tilde{A}$, which can be systematically implemented as a maximum matching problem of a bipartite graph[19].
Lemma 1
[20] Consider the matrix $\tilde{A}$ in eq. (9). There exist row and column permutation matrices $P$ and $Q$ that transform $\tilde{A}$ into the upper-triangular block matrix by
${P}\tilde{A}{Q}=\begin{bmatrix}\tilde{A}_{u}&\tilde{A}_{ud}\\ O&\tilde{A}_{d}\end{bmatrix},$ (10)
where $\tilde{A}_{d}\in\mathbb{R}^{N_{d}\times N_{d}}$ is a block-triangular matrix with $\eta$ diagonal blocks satisfying $\text{rank}(\tilde{A}_{d})=N_{d}$, and $\tilde{A}_{u}\in\mathbb{R}^{(M-N_{d})\times(N-N_{d})}$ is a matrix satisfying $\text{rank}(\tilde{A}_{u})=M-N_{d}$. Among such transformations, the partition maximizing $\eta$ is unique, and each diagonal block in $\tilde{A}_{d}$ is uniquely determined up to row and column permutations within that block.
It is known that the DM decomposition of eq. (10) produces such block-triangular form that maximally separates the welldetermined component with the finest partitions from the original matrix [20]. This allows for extracting the moments and their derivatives that can be fully determined from the underdetermined moment equations. The next lemma shows that the maximal welldetermined block of $\Phi$ in eq. (9) is obtained using $\tilde{A}_{d}$ in Lemma 1.
Lemma 2
There exist row and column permutations that transform $\Phi$ in eq. (9) into the upper-triangular block matrix
$\Psi=\begin{bmatrix}\tilde{A}_{u1}&\tilde{A}_{u2}&\tilde{A}_{u3}\\ O&\tilde{A}_{d}&\tilde{A}^{k}_{d}\\ O&O&\tilde{A}_{d}\end{bmatrix},$ (11)
where $\tilde{A}_{d}\in\mathbb{R}^{N_{d}\times N_{d}}$ is a block-triangular matrix with $\eta$ diagonal blocks defined in eq. (10), and $\tilde{A}_{d}^{k}\in\mathbb{R}^{N_{d}\times N_{d}},\tilde{A}_{u1}\in\mathbb{R}^{2(M-N_{d})\times 2(N-N_{d})},\tilde{A}_{u2}\in\mathbb{R}^{2(M-N_{d})\times N_{d}}$ and $\tilde{A}_{u3}\in\mathbb{R}^{2(M-N_{d})\times N_{d}}$ are real-valued matrices defined accordingly. Moreover, this decomposition maximizes the diagonal blocks among all row and column permutations of $\Phi$, and each diagonal block of $\Psi$ is unique up to row and column permutations within the block.
The proof of this proposition is shown in Appendix -A. Using Lemma 2, we rearrange eq. (8) to obtain
$\Psi\bm{\nu}_{\Psi}=\bm{b}_{\Psi},$ (12)
where $\bm{\nu}_{\Psi}\in\mathbb{R}^{2N}$ and $\bm{b}_{\Psi}\in\mathbb{R}^{2M}$ are the vectors obtained by rearranging the entries of $\bm{\nu}_{\Phi}$ and $\bm{b}_{\Phi}$ according to the transformation from the matrix $\Phi$ to the matrix $\Psi$. Consequently, all welldetermined equations are extracted with the block-triangular coefficient matrix $\Psi$ as illustrated in Fig. 1, where $\bullet$ represents an arbitrary non-zero element. Thus, by applying backward substitution for the linear equations (12), we can systematically determine the entries of $\bm{\nu}_{\Psi}$ that are zero regardless of the coefficients, from which invariant moments are identified.
Figure 1: Illustration of a matrix transformation that extracts a welldetermined upper-triangular equation (shown in green). In this example, $\mathbb{E}[x_{3}]^{(1)}=0$ holds from the structural information of the equation.
Example: Consider the reaction network with molecular sequestrations $\Gamma$, where the stoichiometry vector $\bm{s}_{j}$ and the propensity function $w_{j}(\bm{x})$ are given in Table I.
TABLE I: Definitions of the chemical reaction network in Example
$R_{j}$ Stoichiometry vector $\bm{s}_{j}$ Propensity function $w_{j}(\bm{x})$
$R_{1}$ $[1,0,0]^{\top}$ $\theta_{1}$
$R_{2}$ $[0,1,0]^{\top}$ $\theta_{2}x_{1}$
$R_{3}$ $[0,0,1]^{\top}$ $\theta_{3}x_{2}$
$R_{4}$ $[-1,0,-1]^{\top}$ $\theta_{4}x_{1}z_{3}$
$R_{5}$ $[0,-2,0]^{\top}$ ${\theta_{5}x_{2}(x_{2}-1)}/2$
The truncated stationary moment equations up to the second order moments is then given by
$\begin{bmatrix}0&0&0&-\theta_{4}\\ \theta_{2}&\theta_{5}&-\theta_{5}&0\\ 0&\theta_{3}&0&-\theta_{4}\\ \end{bmatrix}\begin{bmatrix}\mathbb{E}[x_{1}]\\ \mathbb{E}[x_{2}]\\ \mathbb{E}[x_{2}^{2}]\\ \mathbb{E}[x_{1}x_{3}]\end{bmatrix}=-\begin{bmatrix}\theta_{1}\\ 0\\ 0\\ \end{bmatrix}.$ (13)
Here, our goal is to find invariant moments subject to the perturbation to the parameter $\theta_{3}$ of the reaction $R_{3}$. To this end, we formulate the linear equations (7) from eq. (13) and apply the DM decomposition to obtain eq. (12), or
$\begin{bmatrix}\bullet&\bullet&0&0&\bullet&0&0&0\\ 0&0&\bullet&\bullet&0&0&\bullet&0\\ 0&0&0&0&\bullet&\bullet&\bullet&0\\ 0&0&0&0&0&\bullet&0&0\\ 0&0&0&0&0&0&\bullet&\bullet\\ 0&0&0&0&0&0&0&\bullet\end{bmatrix}\begin{bmatrix}\mathbb{E}[x_{2}^{2}]^{(1)}\\ \mathbb{E}[x_{1}]^{(1)}\\ \mathbb{E}[x_{2}^{2}]\\ \mathbb{E}[x_{1}]\\ \mathbb{E}[x_{2}]^{(1)}\\ \mathbb{E}[x_{1}x_{3}]^{(1)}\\ \mathbb{E}[x_{2}]\\ \mathbb{E}[x_{1}x_{3}]\end{bmatrix}=\begin{bmatrix}0\\ 0\\ 0\\ 0\\ 0\\ \bullet\end{bmatrix}.$ (14)
The resulting coefficient matrix becomes upper-triangular, and thus, by applying the backward substitution for linear equations, we can verify that ${\mathbb{E}}[x_{1}x_{3}]^{(1)}=0$ regardless of the parameter value $\bm{\theta}$.
In the next subsection, we consider a graph-based representation of eq. (12) and propose a more systematic procedure to analyze the invariance of moments with respect to parameters.
III-C Graph-based identification of invariant moments
We generalize the idea of the backward substitution shown in Example to systematically identify the invariant moments. To this end, we define $\mathcal{R}_{i}$ and $\mathcal{C}_{i}$ as the set of rows and columns corresponding to the $i$-th diagonal block of the matrix $\Psi$ ($i=0,1,2,\ldots,2\eta$), where the subscript 0 designates the upper-left block $\tilde{A}_{u1}$ of $\Psi$. Corresponding to this partition, we define the $12\eta+$ 1 partitions of $\bm{\nu}_{\Phi}$ and $\bm{b}_{\Psi}$ as
${\bm{\nu}}_{\Psi}=[\tilde{\bm{\nu}}_{0},\tilde{\bm{\nu}}_{1},\dots,\tilde{\bm{\nu}}_{2\eta}]^{\top},\>{\bm{b}}_{\Psi}=[\tilde{\bm{b}}_{0},\tilde{\bm{b}}_{1},\dots,\tilde{\bm{b}}_{2\eta}]^{\top},$ (15)
where $\tilde{\bm{\nu}}_{i+\eta}$ and $\tilde{\bm{\nu}}_{i}\>(i=1,2,\dots,\eta)$ represent the stationary moments and those partial derivative, respectively. In what follows, we interpret these variables as the nodes in a graph $G(\mathcal{V},\mathcal{E})$ with node set $\mathcal{V}$ and edge set $\mathcal{E}$, and we analyze the parameter invariance of the stationary moments through the reachability properteies of the graph. More formally, let $V_{i}$ denote the node corresponding to the variables $\tilde{\bm{\nu}}_{i}$, and we define the set of nodes by $\mathcal{V}:=\{V_{1},V_{2},\ldots,V_{2\eta}\}$. An edge $(V_{q},V_{p})$ is introduced from $V_{q}$ to $V_{p}$ if the corresponding block in $\Psi$ is non-zero, i.e, $\Psi[\mathcal{R}_{p},\mathcal{C}_{q}]\neq O$, where $\Psi[\mathcal{R}_{p},\mathcal{C}_{q}]$ is the $(p,q)$ block of $\Psi$ associated with the row set $\mathcal{R}_{p}$ and the column set $\mathcal{C}_{q}$. Thus, the set of all edges is defined by
:=$\mathcal{E}:=\{(V_{q},V_{p})\in\mathcal{V}\times\mathcal{V}\mid\Psi[\mathcal{R}_{p},\mathcal{C}_{q}]\neq O\}.$ (16)
The following theorem shows a graph-based analysis method for identifying the invariant moments.
Theorem 1
Consider the graph $G(\mathcal{V},\mathcal{E})$ induced by eq. (12). Define a set of source nodes by $\mathcal{S}:=\{V_{i}\in\mathcal{V}\mid\tilde{\bm{b}}_{i}\neq\bm{0}\}(\subseteq\mathcal{V})$. Then, $\tilde{\bm{\nu}}_{i}=\bm{0}$ holds for every equation (12) if $V_{i}$ is unreachable from all source nodes $\mathcal{S}$. $\tilde{\bm{\nu}}^{(1)}_{r}=\bm{0}$.
Proof:
In general, the equations for the row set $\mathcal{R}_{i}$ are given by
$\Psi[\mathcal{R}_{i},\mathcal{C}_{i}]\tilde{\bm{\nu}}_{i}+\sum_{q\in\mathcal{N}_{i}}\Psi[\mathcal{R}_{i},\mathcal{C}_{q}]\tilde{\bm{\nu}}_{q}=\tilde{\bm{b}}_{i}$ (17)
where $\mathcal{N}_{i}:=\{j\mid(V_{j},V_{i})\in\mathcal{E}\}$. This implies that $\tilde{\bm{\nu}}_{i}=\bm{0}$ if the variables $\tilde{\bm{\nu}}_{q}\ (q\in\mathcal{N}_{i})$ and $\bm{\tilde{b}}_{i}$ in eq. (17) are zero ()($\star$)( ⋆ ). Since $V_{i}$ is unreachable from the source nodes, $\tilde{\bm{b}}_{q}=0$ for every equation associated with an upstream nodes $V_{q}$ along the directed paths leading to $V_{i}$. Moreover, the equations for the most upstream node $V_{j}$ reduces to $\Psi[\mathcal{R}_{j},\mathcal{C}_{j}]\tilde{\bm{\nu}}_{j}=\bm{0}$, which implies $\tilde{\bm{\nu}}_{j}=\bm{0}$ as $\Psi[\mathcal{R}_{j},\mathcal{C}_{j}]$ is a welldetermined block. Thus, by recursively applying ()($\star$)( ⋆ ) from the most upstream node to $V_{i}$, we have $\tilde{\bm{\nu}}_{i}=0$. ∎
Theorem 1 represents the idea of backward substitution in Example using the concept of unreachability in directed graphs. Since efficient algorithms for determining reachability in directed graphs have been extensively studied [22, 23, 24], it is possible to systematically and efficiently identify invariant stationary moments subject to the perturbation of the parameter $\theta_{k}$. More specifically, $\mathcal{M}_{k}$ can be obtained by applying Theorem 1 to $\tilde{\bm{\nu}}_{i}\>(i=1,2,\dots,\eta)$, by which we can identify the stationary moments whose partial derivatives are structurally determined to be zero.
IV Application example
Consider the chemical reaction network in Fig. 2(a), where the network consists of $n=6$ molecular species and $m=11$ reactions. Here, we aim to identify the invariant stationary moments under the variation of $\theta_{3}$, i.e., $\mathcal{M}_{3}$ in eq. (3).
For this purpose, we first formulate the stationary moment equation (5) up to the second order moments, which results in 27 equations with 32 variables. Next, we apply the DM decomposition to the structure matrix $\tilde{A}$ to extract the square matrix $\tilde{A}_{d}$ (Lemma 1). The resulting matrix is shown in Fig. 2(b), which indeed becomes the upper-triangular form. In particular, the square matrix $\tilde{A}_{d}$ with $N_{d}=7$ dimensions with $\eta=7$ diagonal blocks (scalars) is successfully extracted from the underdetermined moment equation. Applying the same row and column permutations, i.e., $P$ and $Q$ in Lemma 1, to eq. (8), we further obtain $\Psi\bm{\nu}_{\Psi}=\bm{b}_{\Psi}$ in eq. (12) (Lemma 2 and Appendix -A). Figure 2(c) shows the welldetermined part of the equation $\Psi\bm{\nu}_{\Psi}=\bm{b}_{\Psi}$, which becomes $142N_{d}=$ 14 dimensions in the upper-triangular form. Finally, using Theorem 1, we explore the set of moments $\mathcal{M}_{3}$. Specifically, a graph $G(\mathcal{V},\mathcal{E})$ with 14 nodes with two source nodes at $V_{8}$ and $V_{14}$ is defined since $b_{8}\neq 0$ and $b_{14}\neq 0$. Then, the reachable states are examined by tracing all reachable nodes from the sources (see the orange lines in Fig. 2(c)).
Figure 2: Application example. (a) a network of feedback reactions, (b) DM decomposition of $\tilde{A}$, and (c) identification of invariant moments based on the backward substitution of eq. (12)
From these results, it is concluded that $\mathbb{E}[x_{1}x_{6}],\mathbb{E}[x_{5}],\mathbb{E}[x_{4}],$ and $\mathbb{E}[x_{3}]$ are invariant moments $\mathcal{M}_{3}$ with respect to the variation of $\theta_{3}$. These results demonstrate the systematic and computationally efficient feature of the proposed approach, which leverages the DM decomposition of the coefficient matrix along with the unreachability analysis of the graph. The identified invariant moments provide deeper insights into homeostatic behaviors of biomolecular systems.
V Conclusion
In this paper, we have proposed a systematic method for analyzing the parameter invariance of stationary moments in general chemical reaction networks by leveraging the non-zero structure of the moment equations. First, by applying the Dulmage-Mendelsohn (DM) decomposition to the moment equations, we have extracted welldetermined subequations from the underdetermined equations and have revealed the hierarchical structure among stationary moments. Furthermore, by performing unreachability analysis on a directed graph based on this structure, we have demonstrated that stationary moments whose partial derivatives with respect to parameters are structurally zero can be systematically identified. The proposed method enables rigorous parameter invariance analysis for a broad class of stochastic chemical systems, offering new insights into the homeostasis of chemical reaction networks.